OpenStudy (anonymous):

Find four consecutive even integers whose sum is 244.

OpenStudy (anonymous):

so I did this, and couldn't get 244 with integers, is there something I DON'T know?

OpenStudy (anonymous):

it cant be 58 60 62 and 64, because I've already been told that is wrong

OpenStudy (texaschic101):

four consecutive even integers... x, x + 2, x + 4, x + 6 now lets make an equation... (x) + (x + 2) + (x + 4) + (x + 6) = 244 (combine like terms) 4x + 12 = 244 4x = 244 - 12 4x = 232 x = 232/4 x = 58 x + 2 = 58 + 2 = 60 x + 4 = 58 + 4 = 62 x + 6 = 58 + 6 = 64 Your numbers are 58,60,62,64 That is not wrong. You got it right.